OpenStudy (anonymous):

The equation of circle having a diameter with endpoints (-3, 1) and (5, 7) is

OpenStudy (anonymous):

The answers available are (x - 4)² + (y - 3)² = 25 (x - 1)² + (y - 4)² = 25 (x - 1)² + (y - 4)² = 100

hartnn (hartnn):

distance between end points of diameter will be twice the radius, so first find the distance between (-3, 1) and (5, 7)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

OpenStudy (anonymous):

So (-3-5)^2 + (1-7)^2

hartnn (hartnn):

yes, simplify and take square root..

OpenStudy (anonymous):

But I don't understand how I'd end up with one o these as an answer (x - 4)² + (y - 3)² = 25 (x - 1)² + (y - 4)² = 25 (x - 1)² + (y - 4)² = 100

OpenStudy (anonymous):

*of

hartnn (hartnn):

when you get the radius, just square it. for radius you will first need diameter...

OpenStudy (anonymous):

I'm confused.d..q.wdpfqwjfwpefv.:(

hartnn (hartnn):

(-3-5)^2 + (1-7)^2 = 8^2 +6^2 = 100 its square root will be 10 radius = 5 to get the centre, find the midpoint of (-3, 1) and (5, 7) can u ?

OpenStudy (anonymous):

I dontknow how to do any of it.. I feel sooo dumb.

OpenStudy (anonymous):

don't know*

hartnn (hartnn):

mid-point we did recently, right ? just find the average of co-ordinates...

hartnn (hartnn):

(-3+5)/2 =... ? (1+7)/2 =... ?

OpenStudy (anonymous):

0 4

hartnn (hartnn):

actually (1,4) so, the equation of circle will be \((x-h)^2+(y-k)^2=r^2 \\ (x-1)^2+(y-4)^2 = 5^2\)

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

So it'd be (x - 4)² + (y - 3)² = 25

OpenStudy (anonymous):

?

hartnn (hartnn):

no.... (x - 1)² + (y - 4)² = 25